OpenStudy (anonymous):

Find the period, range, and amplitude of the cosine function. y=-4 cos 8x

OpenStudy (anonymous):

@jim_thompson5910 @Directrix @dan815 help?

jimthompson5910 (jim_thompson5910):

what do you have so far?

OpenStudy (anonymous):

Nothing?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

in general, if we have y = A*cos(Bx-C) + D we know that |A| is the amplitude T = 2pi/B is the period C/B is the phase shift y = D is the midline

jimthompson5910 (jim_thompson5910):

in this problem, what is the amplitude?

OpenStudy (anonymous):

I'm sorry, I suck at this. How do you get the amplitude?

jimthompson5910 (jim_thompson5910):

compare y = A*cos(Bx-C) + D to y=-4*cos(8x)

jimthompson5910 (jim_thompson5910):

notice how A matches with the -4

jimthompson5910 (jim_thompson5910):

|A| is the amplitude

OpenStudy (anonymous):

So the amplitude is -4?

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

|A| = |-4| = ??

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

Okay, the amplitude is 4, is the period 1/4 or 1/4 pi?

jimthompson5910 (jim_thompson5910):

T = 2pi/B is the period in this case, B = 8, so T = 2pi/B T = 2pi/8 T = pi/4

jimthompson5910 (jim_thompson5910):

pi/4 is the period and that's the same as (1/4)pi pi/4 is probably more common notation

jimthompson5910 (jim_thompson5910):

what is the midline?